### v1

@@ 2.1 UNet-base Encoder-Decoder
UNet is a highly successful foundational network architecture in object detection. It utilizes an encoder-decoder structure to learn semantic information at different levels. Qin et al.[15] further improved detection performance by proposing a nested structure that stacks multiple UNet models on top of each other (as shown in Fig.(1)). The network consists of multiple RSU (Recursive Skip U-Net) modules, which can be seen as miniaturized UNet models. The input to each RSU module comes from different levels of the encoding features. Through this multi-scale feature fusion, the network can learn rich contextual information. Compared to UNet, UNet has a deeper hierarchy, which expands the receptive field. Additionally, the nested U-connection enables effective feature fusion from different levels.
@@ 4.1 Evaluating indicator
where P and P represent the heavy rainfall amount predicted by the model and observed by radar, respectively. HSS is a statistical method used to measure the accuracy of predictions and evaluate the performance of classifiers in binary classification problems. The HSS ranges from -1 to 1, where 1, 0, and -1 denote accurate, random, and completely incorrect predictions. FSS is defined as Eq.(4),respectively:
@@ 4.1.1 SRNDiff Implementation Details
We trained two different models: SRNDiff (activating the attention block in the conditional network when the feature map size is less than or equal to 32), and the SRNDiff model without the attention block. The diffusion steps were set to 1000, and the Adam Optimizer was used with a learning rate of . Each model underwent approximately 33 million training steps and was trained for 72 hours on 10 A6000 GPUs. For additional training configuration details of other models, please refer to Table 1.
model |
(Gen)optimize |
Discriminator optimize |
step |
BatchSize |
UNet-GAN |
Adam |
Adam |
20M |
8 |
UNet-GAN* |
Adam |
Adam |
20M |
8 |
UNet-GAN |
Adam |
Adam |
20M |
8 |
UNet-GAN* |
Adam |
Adam |
20M |
8 |
DGMR* |
Adam |
Adam |
20M |
8 |
DGMR |
Adam |
Adam |
20M |
8 |
SRNDiff |
Adam |
33M |
32 | |
SRNDiff |
Adam |
33M |
32 |
@@ 4.1.3 Result
model |
UNet-GAN |
0.94 |
0.38 |
0.29 |
0.17 |
UNet-GAN* |
0.94 |
0.41 |
0.32 |
0.19 |
UNet-GAN |
0.95 |
0.38 |
0.30 |
0.16 |
UNet-GAN* |
0.95 |
0.29 |
0.18 |
0.08 |
DGMR |
0.91 |
0.33 |
0.25 |
0.16 |
DGMR* |
0.94 |
0.36 |
0.34 |
0.16 |
SRNDiff |
0.96 |
0.48 |
0.329 |
0.20 |
SRNDiff |
0.96 | 0.49 |
0.334 |
0.20 |
model |
UNet-GAN |
0.25 |
0.20 |
0.13 |
UNet-GAN* |
0.27 |
0.23 |
0.16 |
UNet-GAN |
0.24 |
0.14 |
0.07 |
UNet-GAN* |
0.26 |
0.22 |
0.13 |
DGMR |
0.23 |
0.19 |
0.13 |
DGMR* |
0.24 |
0.22 |
0.13 |
SRNDiff |
0.31 |
0.24 |
0.16 |
SRNDiff |
0.32 | 0.25 | 0.17 |
We use the CSI index to analyze the performance of UNet-GAN, DGMR, UNet-GAN,SRNDiff and SRNDiff in predicting ground rainfall for different rainfall intensities. The Table (2) presents the prediction accuracies of these models under different precipitation intensities. When the precipitation intensity is less than , the prediction accuracies of various methods are close, with the diffusion model slightly outperforming the GANs-based method. However, with increasing rainfall, the advantages of the diffusion model gradually become evident. Specifically, when the rainfall is greater than , the SRNDiff accuracy outperforms UNet-GAN by up to 8 percentage points in prediction accuracy. Even at precipitation intensities greater than 4 mm/h and 8 mm/h, the diffusion model maintains a certain accuracy advantage. The HSS scores of these models are presented in Table 2. Under rainfall intensity greater than 2 mm/h, SRNDiff outperforms GANs-based prediction methods by 4 percentage points. Moreover, SRNDiff and SRNDiff also demonstrate good performance in predicting rainfall exceeding 4 mm/h and 8 mm/h. Among them, SRNDiff achieves the highest HSS score, followed by SRNDiff with the second highest. These two evaluation metrics are assessed on a per-pixel basis, thus neither reflects spatial accuracy nor distribution similarity.
model |
FSS |
MSE |
UNet-GAN |
0.53 |
2.90 |
UNet-GAN* |
0.69 |
1.40 |
UNet-GAN |
0.59 |
2.54 |
UNet-GAN* |
0.59 |
1.13 |
DGMR |
0.42 |
2.32 |
DGMR* |
0.63 |
1.88 |
SRNDiff |
0.77 | 0.66 |
SRNDiff |
0.77 | 0.66 |
@@
The visual analysis shows that the shapes of most generated images predicted by the GANs models deviate from the ground truth. These models produce images that prioritize large rainfall areas while overlooking smaller rainfall areas or isolated pixels, leading to overpredicting high rainfall amounts in the generated images. Although the improved loss function in the DGMR* , UNet-GAN* and UNet-GAN* model enhances the accuracy of the generated images compared to DGMR and UNet-GAN, inaccuracies in predicting large rainfall areas exist, particularly for UNet-GAN*. Furthermore, these models lose some edge detail information in their predictions.In comparison, SRNDiff and SRNDiff demonstrate shapes closer to the ground truth and perform better in predicting heavy rainfall over a large area. The calculated evaluation metrics confirm that SRNDiff outperforms all other models.
@@ 4.2 Conclusion
This study comprehensively analyzes the performance of GAN-based and diffusion-based models in precipitation prediction. The results show significant improvements in CSI and HSS metrics for SRNDiff and SRNDiff, especially in scenarios with moderate to heavy rainfall. They outperform GANs-based methods by almost 6 percentage points in FSS and MSE metrics. SRNDiff performs better than SRNDiff in terms of CSI and HSS metrics, while the other metrics are relatively close.
The exceptional performance of SRNDiff can be attributed to two key factors. First, the use of the diffusion model provides stable training and high-quality generation. Second, SRNDiff incorporates a UNet network based on nested shapes and possesses a flexible attention activation mechanism, which efficiently extracts deep features for image encoding and decoding. Additionally, the designed end-to-end prediction model enables joint optimization of the encoder and denoising network, allowing the encoder’s learned features to better adapt to the denoising network’s requirements. Compared to independent training, the end-to-end mechanism allows all modules of the network to collaborate and jointly accomplish the task of predicting high-resolution rainfall images.

### v2

@@ 2.1 UNet-base Encoder-Decoder
UNet is a highly successful foundational network architecture in object detection. It utilizes an encoder-decoder structure to learn semantic information at different levels. Qin et al.[15] further improved detection performance by proposing a nested structure that stacks multiple UNet models on top of each other (as shown in Fig.(1)). The network consists of multiple RSU (Recursive Skip U-Net) modules, which can be seen as miniaturized UNet models. The input to each RSU module comes from different levels of the encoding features. Through this multi-scale feature fusion, the network can learn rich contextual information. Compared to UNet, U2Net has a deeper hierarchy, which expands the receptive field. Additionally, the nested U-connection enables effective feature fusion from different levels.
@@ 4.1 Evaluating indicator
where Pfi and Poi represent the heavy rainfall amount predicted by the model and observed by radar, respectively. HSS is a statistical method used to measure the accuracy of predictions and evaluate the performance of classifiers in binary classification problems. The HSS ranges from -1 to 1, where 1, 0, and -1 denote accurate, random, and completely incorrect predictions. FSS is defined as Eq.(4),respectively:
@@ 4.1.1 SRNDiff Implementation Details
We trained two different models: SRNDiffatten (activating the attention block in the conditional network when the feature map size is less than or equal to 32), and the SRNDiff model without the attention block. The diffusion steps were set to 1000, and the Adam Optimizer was used with a learning rate of . Each model underwent approximately 33 million training steps and was trained for 72 hours on 10 A6000 GPUs. For additional training configuration details of other models, please refer to Table 1.
model | (Gen)optimize | Discriminator optimize | step | BatchSize |
U2Net-GAN | Adam | Adam | 20M | 8 |
U2Net-GAN* | Adam | Adam | 20M | 8 |
UNet-GAN | Adam | Adam | 20M | 8 |
UNet-GAN* | Adam | Adam | 20M | 8 |
DGMR* | Adam | Adam | 20M | 8 |
DGMR | Adam | Adam | 20M | 8 |
SRNDiff | Adam | 33M | 32 | |
SRNDiffatten | Adam | 33M | 32 |
@@ 4.1.3 Result
model | |||||
U2Net-GAN | 0.94 | 0.38 | 0.29 | 0.17 | |
U2Net-GAN* | 0.94 | 0.41 | 0.32 | 0.19 | |
UNet-GAN | 0.95 | 0.38 | 0.30 | 0.16 | |
UNet-GAN* | 0.95 | 0.29 | 0.18 | 0.08 | |
DGMR | 0.91 | 0.33 | 0.25 | 0.16 | |
DGMR* | 0.94 | 0.36 | 0.34 | 0.16 | |
SRNDiff | 0.96 | 0.48 | 0.329 | 0.20 | |
SRNDiffatten | 0.96 | 0.49 | 0.334 | 0.20 |
model | |||
U2Net-GAN | 0.25 | 0.20 | 0.13 |
U2Net-GAN* | 0.27 | 0.23 | 0.16 |
UNet-GAN | 0.24 | 0.14 | 0.07 |
UNet-GAN* | 0.26 | 0.22 | 0.13 |
DGMR | 0.23 | 0.19 | 0.13 |
DGMR* | 0.24 | 0.22 | 0.13 |
SRNDiff | 0.31 | 0.24 | 0.16 |
SRNDiffatten | 0.32 | 0.25 | 0.17 |
We use the CSI index to analyze the performance of U2Net-GAN, DGMR, UNet-GAN,SRNDiff and SRNDiff atten in predicting ground rainfall for different rainfall intensities. The Table (2) presents the prediction accuracies of these models under different precipitation intensities. When the precipitation intensity is less than , the prediction accuracies of various methods are close, with the diffusion model slightly outperforming the GANs-based method. However, with increasing rainfall, the advantages of the diffusion model gradually become evident. Specifically, when the rainfall is greater than , the SRNDiff accuracy outperforms U2Net-GAN by up to 8 percentage points in prediction accuracy. Even at precipitation intensities greater than 4 mm/h and 8 mm/h, the diffusion model maintains a certain accuracy advantage. The HSS scores of these models are presented in Table 2. Under rainfall intensity greater than 2 mm/h, SRNDiff outperforms GANs-based prediction methods by 4 percentage points. Moreover, SRNDiff and SRNDiffatten also demonstrate good performance in predicting rainfall exceeding 4 mm/h and 8 mm/h. Among them, SRNDiff achieves the highest HSS score, followed by SRNDiff with the second highest. These two evaluation metrics are assessed on a per-pixel basis, thus neither reflects spatial accuracy nor distribution similarity.
model | FSS | MSE |
U2Net-GAN | 0.53 | 2.90 |
U2Net-GAN* | 0.69 | 1.40 |
UNet-GAN | 0.59 | 2.54 |
UNet-GAN* | 0.59 | 1.13 |
DGMR | 0.42 | 2.32 |
DGMR* | 0.63 | 1.88 |
SRNDiff | 0.77 | 0.66 |
SRNDiffAttention | 0.77 | 0.66 |
@@
The visual analysis shows that the shapes of most generated images predicted by the GANs models deviate from the ground truth. These models produce images that prioritize large rainfall areas while overlooking smaller rainfall areas or isolated pixels, leading to overpredicting high rainfall amounts in the generated images. Although the improved loss function in the DGMR* , UNet-GAN* and U2Net-GAN* model enhances the accuracy of the generated images compared to DGMR and UNet-GAN, inaccuracies in predicting large rainfall areas exist, particularly for UNet-GAN*. Furthermore, these models lose some edge detail information in their predictions.In comparison, SRNDiff and SRNDiffatten demonstrate shapes closer to the ground truth and perform better in predicting heavy rainfall over a large area. The calculated evaluation metrics confirm that SRNDiffatten outperforms all other models.
@@ 4.2 Conclusion
This study comprehensively analyzes the performance of GAN-based and diffusion-based models in precipitation prediction. The results show significant improvements in CSI and HSS metrics for SRNDiff and SRNDiffatten, especially in scenarios with moderate to heavy rainfall. They outperform GANs-based methods by almost 6 percentage points in FSS and MSE metrics. SRNDiffatten performs better than SRNDiff in terms of CSI and HSS metrics, while the other metrics are relatively close.
The exceptional performance of SRNDiffatten can be attributed to two key factors. First, the use of the diffusion model provides stable training and high-quality generation. Second, SRNDiffatten incorporates a UNet network based on nested shapes and possesses a flexible attention activation mechanism, which efficiently extracts deep features for image encoding and decoding. Additionally, the designed end-to-end prediction model enables joint optimization of the encoder and denoising network, allowing the encoder’s learned features to better adapt to the denoising network’s requirements. Compared to independent training, the end-to-end mechanism allows all modules of the network to collaborate and jointly accomplish the task of predicting high-resolution rainfall images.